ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 211, 2016

Adiprasito, K. A. and A. Padrol A universality theorem for projectively unique polytopes and a conjecture of Shephard

239

Albuquerque, N., F. Bayart, D. Pellegrino and J. B. Seoane-Sepúlveda Optimal Hardy–Littlewood type inequalities for polynomials and multilinear operators

197

Austin, T. On the failure of concentration for the ℓ_{∞}-ball  221

Böttcher, J., J. Hladký, D. Piguet and A. Taraz An approximate version of the Tree Packing Conjecture

391

Bayart, F. See N. Albuquerque Charette, F. and O. Cornea Categorification of Seidel’s represen-

tation

67

Ciobanu, L., S. Hermiller, D. Holt and S. Rees Conjugacy languages in groups

311

Cornea, O. See F. Charette

de la Salle, M. Towards strong Banach property (T) for SL(3, ℝ)  105

Detomi, E. and A. Lucchini Invariable generation of prosoluble groups

481

Foryś, M., W. Huang, J. Li and P. Oprocha Invariant scrambled sets, uniform rigidity and weak mixing

447

Glebov, R., D. Král’ and J. Volec A problem of Erdős and Sós on 3-graphs

349

Hanke, T., D. Neftin and A. Wadsworth Galois subfields of tame division algebras

367

Hermiller, S. See L. Ciobanu Hladký, J. See J. Böttcher Holt, D. See L. Ciobanu Huang, W. See M. Foryś

Iyama, O. and B. Lerner Tilting bundles on orders on ℙ^{d}  147

Keller, T. M. and Y. Yang Abelian quotients and orbit sizes of solvable linear groups

23

Král’, D. See R. Glebov Lascar, B. and R. Lascar L’inégalité de Melin–Hörmander en car-

actéristiques multiples 13 Lascar, B. and R. Lascar Une condition suffisante pour que le

problème de Cauchy soit bien posé

1

Lascar, R. See B. Lascar Lefèvre, P., Ṁatheron and A. Primot Smoothness, asymptotic

smoothness and the Blum–Hanson property

271

Lerner, B. See O. Iyama Li, J. See M. Foryś Lucchini, A. See E. Detomi Matheron, ˙ See P. Lefèvre Neftin, D. See T. Hanke Oprocha, P. See M. Foryś Padrol, A. See K. A. Adiprasito Pellegrino, D. See N. Albuquerque Piguet, D. See J. Böttcher

Prasad, D. A character relationship on GL_{n}(ℂ)  257

Primot, A. See P. Lefèvre Rees, S. See L. Ciobanu Schein, M. M. and C. Voll Normal zeta functions of the Heisenberg

groups over number rings II — the non-split case

171

Seoane-Sepúlveda, J. B. See N. Albuquerque

Spinas, O. Silver trees and Cohen reals  473

Taraz, A. See J. Böttcher Volec, J. See R. Glebov Voll, C. See M. M. Schein Wadsworth, A. See T. Hanke Yang, Y. See T. M. Keller

Zhang, Z. A matrix description for K_{1} of graded rings  45

Zywina, D. Abelian varieties over large algebraic fields with infinite torsion 493